If the perimeter of a fence is 76 feet what are the dimensions of the length and width if the length is (x+4) and width is (2x-23).
Answer by addingup(3677)

2Length+2Width = Perimeter
L = (x+4)
W = (2x-23)
---------------------
2(x+4)+2(2x-23) = 76
Distributive property of multiplication:
2x+8+4x-46 = 76
consolidate like terms:
2x+4x+8-46 = 76
6x-38 = 76
add 38 on both sides
6x = 114
x = 19
----------------------
L = 19+4 = 23
W = 2(19)-23 = 38-23 = 15
------------------------------
check:
2(23)+2(15) = 76
46+30 = 76 Correct
:
John